OpenStudy (anonymous):

Two numbers are 10 units away in different directions from their midpoint, m, on a number line. The product of the numbers is –99. Which equation can be used to find m, the midpoint of the two numbers?

OpenStudy (anonymous):

@acxbox22

OpenStudy (anonymous):

is it (x-10)(x+10)=-99?

jimthompson5910 (jim_thompson5910):

"Two numbers are 10 units away in different directions from their midpoint, m, on a number line" so these two numbers are 20 units apart (2*10 = 20) if x and y are the two numbers, then x-y = 20 where x is the larger number we also know that x*y = -99 so one of the numbers is positive while the other is negative

jimthompson5910 (jim_thompson5910):

x-y = 20 x-y+y = 20+y y+20 = x y = x - 20 ------------- x*y = -99 x*(x-20) = -99 solve for x and use this to find y

OpenStudy (anonymous):

how do i find the equation out of that?

OpenStudy (anonymous):

(m – 5)(m + 5) = 99 (m – 10)(m + 10) = 99 m2 – 25 = –99 m2 – 100 = –99 these are my options. i thought it was b but then i realized it has to equal -99

jimthompson5910 (jim_thompson5910):

oh they want it in terms of m

jimthompson5910 (jim_thompson5910):

let me think

jimthompson5910 (jim_thompson5910):

ok if m is our reference point, then the first number is m+10 (since you count 10 units above the midpoint) and the second number is m-10 agreed so far?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

do yo multiply them together?

jimthompson5910 (jim_thompson5910):

so (m-10)*(m+10) is the product of the two numbers we can FOIL that out to get m^2 - 10m + 10m - 100 = m^2 - 100

jimthompson5910 (jim_thompson5910):

or you can use the difference of squares rule as a shortcut

OpenStudy (anonymous):

okay

jimthompson5910 (jim_thompson5910):

so that's why the answer is m^2 - 100 = -99

OpenStudy (anonymous):

i understand

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

you're welcome